Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-044_2000-050

Mission Operations Days: 2000/044 to 2000/050
February 13, 2000 00:00:00 GMT to February 19, 2000 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
MODIS has no known Anomalies
ComponentState(s)Comment
BlackbodyA On (290 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, ClosedNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On, #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA On; B OffNominal
Flight SoftwareRev BDNominal
Inhibit Ids SetNoneNominal
TMONs EnabledNoneNominal

This Week’s Completed MODIS Activities:

Saturday, February 12th, 2000

2000/043 20:43 - 150s orbit attainment burn executed nominally
2000/043 20:43:27 - SSR Disabled due to thruster burn. Data loss begins here.
2000/043 22:15:29 - SSR Re-enabled

Sunday, February 13th, 2000

2000/044 15:00 - started controlling on the focal planes
2000/044 16:51 - MODIS turn off
2000/044 16:55 - PS2 Off
2000/044 16:59:46 - MODIS turn on PS1, CPA
2000/044 17:02:28 - MODIS transition to Science Mode, A-side
2000/044 17:06:03 - Set formatter to night mode
2000/044 18:18:59 - Begin SDSM Functional Test
2000/044 18:27:35 - End SDSM Functional Test
2000/044 18:30 - Clear inhibit Ids 31, 32, 33, 53, 58
2000/044 20:43:27 – 21:01:15 ATC - - SDSM Screened double collect (OA16) / EDS
2000/044 21:15:00 – 21:19:31 ATC - - PC Ecal (OA28)
2000/044 21:20:00 – 21:22:31 ATC - - PV Ecal (OA27)
2000/044 21:50:00 Real-time - - Set BB duty cycle to 100%
2000/044 22:21:43 – 22:39:36 ATC - - SDSM Open double collect (OA15) / EDS

Monday, February 14th, 2000

2000/045 15:42:15 - SSR buffer disabled. Data loss.
2000/045 15:47:50 - SSR buffer reenabled
2000/045 17:00 ATC - - OA26: BB On to 315K at 100% duty cycle
2000/045 20:12:00 - 300s orbit attainment burn
2000/045 22:22:10 - DC Restore (PV and PC) Off
2000/045 22:24:40 - Sector Rotation to –3072
2000/045 22:32:30 - Sector Rotation to 0
2000/045 22:33:45 - DC Restore (PV and PC) On

Tuesday, February 15th, 2000

2000/046 01:00:00 - BB Off
2000/046 14:25:00 - Set BB to 33% duty cycle
2000/046 15:52:55 – 16:20:00 - SRCA 1W Spatial (SRCA Functional)
2000/046 17:00:00 - BB on to 315K at 33% duty cycle for 8 hours

Wednesday, February 16th, 2000

2000/047 01:00 ATC - Blackbody turn off
2000/047 14:30:00 – 15:03:33 ATC - - SRCA Full Radiometric
2000/047 15:38 Real-time - - Turn on BB to 290.15K at 100% duty cycle
2000/047 19:46 - 320s orbit attainment burn; data loss expected
2000/047 21:55:41 Real-time - - Set BB Dcycle to 33%
2000/047 21:56:44 Real-time - - Turn on EDS (quicklook)
2000/047 21:58:00 Real-time - - SMIR Vdet Sweeps, Itwk = 79
2000/047 22:07:59 Real-time - - SMIR Vdet Sweeps, Itwk = 90
2000/047 22:15:10 Real-time - - SMIR Vdet Sweeps, Itwk = 100
2000/047 22:22:13 Real-time - - SMIR Vdet Sweeps, Itwk = 110
2000/047 22:29:19 Real-time - - Reset default values (Itwk = 79)
2000/047 22:30:28 Real-time - - Turn off EDS (quicklook)

Thursday, February 17th, 2000

2000/048 14:43:00 – 15:40:31 ATC - - SRCA Full Spatial
2000/048 19:26:41 Real-time - - Turn on EDS (quicklook)
2000/048 19:28:56 Real-time - - LWIR Vdet Sweeps, Itwk = 79
2000/048 19:36:16 Real-time - - LWIR Vdet Sweeps, Itwk = 85
2000/048 19:44:11 Real-time - - Turn off EDS (quicklook)
*Note that values for LWIR Vdet/Itwk were not set back to nominal between this set and the next set of sweeps.
2000/048 22:41:55 Real-time - - Turn on EDS (quicklook)
2000/048 22:42:57 Real-time - - LWIR Vdet Sweeps, Itwk = 90
2000/048 22:50:17 Real-time - - LWIR Vdet Sweeps, Itwk = 95
2000/048 22:58:01 Real-time - - Reset default values (Itwk = 79)
2000/048 22:59:48 Real-time - - Turn off EDS (quicklook)

Friday, February 18th, 2000

2000/049 13:09:03 – 14:00:46 ATC - - SRCA Full Spectral, 30W Part 1 (macro 18)
2000/049 15:03:25 – 15:46:17 ATC - - SRCA Full Spectral, 30W Part 2 (macro 19)
2000/049 16:25:31 – 17:16:43 ATC - - SRCA Full Spectral, 10W Part 1 (macro 20)
2000/049 18:20:29 – 19:05:42 ATC - - SRCA Full Spectral, 10W Part 2 (macro 21)

Saturday, February 19th, 2000

2000/050 17:16:34 Real-time - - Set SMIR FPA to B side
2000/050 17:17:47 Real-time - - Turn on EDS (quicklook)
2000/050 17:18:51 Real-time - - SMIR-B Vdet Sweeps, Itwk = 79
2000/050 17:26:14 Real-time - - SMIR-B Vdet Sweeps, Itwk = 90
2000/050 17:17:47 Real-time - - Turn off EDS (quicklook)
2000/050 18:55:16 Real-time - - Turn on EDS (quicklook)
2000/050 18:56:16 Real-time - - SMIR-B Vdet Sweeps, Itwk = 100
2000/050 19:03:18 Real-time - - SMIR-B Vdet Sweeps, Itwk = 110
2000/050 19:10:25 Real-time - - Reset SMIR-B default values, (Itwk = 79)
2000/050 19:11:13 Real-time - - Turn off EDS (quicklook)
2000/050 19:12:05 Real-time - - Set SMIR FPA back to A side

This Week's Scheduled MODIS Activities Not Completed:

None.

Upcoming MODIS Events:

Saturday, February 19, 2000

050/20:40 Real-time - - Set BB duty cycle to 100%
050/21:12:22 ATC - - OA16: SD/SDSM Screened

Sunday, February 20th, 2000

2000/051 03:46:55 ATC - - OA16: SD/SDSM Screened
2000/051 08:42:51 ATC - - OA16: SD/SDSM Screened
2000/051 14:00:00 ATC - - Sector shift to –3072, BB at 315K (EDS ~ 5 minutes)
2000/051 14:30 Real-time - - Set BB duty cycle to 33% and temp to 290K
2000/051 15:17:25 ATC - - OA15: SD/SDSM Open
2000/051 20:13:27 ATC - - OA16: SD/SDSM Screened
2000/051 21:14 Real-time - - Turn PV-LW FPA off, Set SMIR FPA to B side
2000/051 21:16 Real-time - - Turn on EDS (quicklook)
2000/051 21:18 Real-time - - SMIR-B Vdet Sweeps, Itwk = 79
2000/051 21:26 Real-time - - SMIR-B Vdet Sweeps, Itwk = 90
2000/051 21:23 Real-time - - SMIR-B Vdet Sweeps, Itwk = 100
2000/051 21:30 Real-time - - SMIR-B Vdet Sweeps, Itwk = 110
2000/051 21:37 Real-time - - Turn off EDS (quicklook)
2000/051 23:00 Real-time - - Turn on EDS (quicklook)
2000/051 23:02 Real-time - - Reset SMIR-B default values, (Itwk = 79)
2000/051 23:10 Real-time - - Turn off EDS (quicklook)
2000/051 23:12 Real-time - - Turn PV-LW FPA back to On, Set SMIR FPA to A side

Monday, February 21st, 2000

2000/052 02:48:27 ATC - - OA16: SD/SDSM Screened
2000/052 09:23:27 ATC - - OA15: SD/SDSM Open
2000/052 10:00 Bit Flip Investigation begins, MODIS OBC Activities suspended.
2000/052 10:18 Real-time - - Enable Formatter Test Data Packets (Bit-Flip Investigation)
2000/052 10:45 Real-time - - Disable Formatter Test Data Packets (Bit-Flip Investigation)
2000/052 12:12 ATC - - Formatter Night -> Day -> Night -> Day -> Night (6 minutes) (Bit-Flip Investigation)

Tuesday, February 22nd, 2000

2000/053 01:00 Bit Flip Investigation ends, MODIS OBC Activities resumed.
2000/053 03:26:32 ATC - - OA16: SD/SDSM Screened, double collection duration
2000/053 08:22:47 ATC - - OA16: SD/SDSM Screened, double collection duration (Sector shift to –3072 and EDS ~18 minutes)
2000/053 13:19:02 ATC - - OA15: SD/SDSM Open, double collection duration
2000/053 ?? Real-time - - Gain table load
2000/053 ?? Real-time - - Ecal Sweeps
2000/053 ?? Real-time - - Ecal Sweeps
2000/053 18:15:17 ATC - - OA16: SD/SDSM Screened, double collection duration
2000/053 23:11:39 ATC - - OA16: SD/SDSM Screened, double collection duration

Wednesday, February 23rd, 2000

2000/054 04:08:13 ATC - - OA15: SD/SDSM Open, double collection duration
2000/054 09:04:47 ATC - - OA16: SD/SDSM Screened, double collection duration
2000/054 14:01:21 ATC - - OA16: SD/SDSM Screened, double collection duration
2000/054 18:57:55 ATC - - OA15: SD/SDSM Open, double collection duration

Thursday, February 24th, 2000

None

Friday, February 25th, 2000

2000/056 ?? Real-time - - Open Nadir Aperture Door (Earth View Data!!)
2000/056 ?? Real-time - - Dump door table

MODIS Anomalies:

There have been two yellow high limits for MOD_VR_PVNIR_RP11V39. SBRS has been notified.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD open. The transition from the B- side to the A-side went well. The functional checkout of the SDSM was nominal. (And there was much rejoicing, again!)

MODIS is in Science Mode on the A-side with the SVD open and BB stable at 290.4K.

The functional checkouts to date: The functional checkout of the SRCA went nominally. The one-watt lamp, source wheel, SIS heater and IR source all reacted as expected. The SRCA’s Spatial, Spectral and Radiometric tests all appear to have gone off without a hitch. All mechanisms and lamps (non-pristine) have been exercised except the Nadir Aperture Door!

MODIS Telemetry Trends:

Telemetry is nominal.

Non-MODIS Significant Events:

Investigation into bit flips continues. No MODIS OBC activities will be planned for primary shift on Monday, February 21st to support bit flip investigation. MODIS has been asked to perform 27 minute Day Rate of Formatter Test Packets and a rapid (4 transitions with 2 minute separations) Formatter Day/Night Rate Toggling. Data will be dumped during this data via various methods. MODIS has been asked to evaluate these varieties of data to determine if bit flips are present in the science data.

Burn Status:

Due to slightly greater than expected overall performance of the previous burns as a result, FDS “had to move and resize burns 7 and 8 in order to achieve the ground track and frozen orbit requirements.” (Note the day change of the final burn in particular to Thursday, February, 23.)

Burn | Start Time | End Time | Duration(sec)

1 | Feb 10 2000 21:15:43.030 | Feb 10 2000 21:16:43.030 | 60.000 (complete)
2 | Feb 12 2000 20:43:22.876 | Feb 12 2000 20:45:52.876 | 150.000 (complete)
3 | Feb 14 2000 20:12:16.920 | Feb 14 2000 20:17:16.920 | 300.000 (complete)
4 | Feb 16 2000 19:45:32.361 | Feb 16 2000 19:50:52.361 | 320.000 (complete)
5 | Feb 18 2000 21:52:11.699 | Feb 18 2000 21:57:31.699 | 320.000 (complete)
6 | Feb 20 2000 19:03:07.853 | Feb 20 2000 19:08:27.853 | 320.000
7 | Feb 22 2000 18:00:-?-.-?-- | Feb 22 2000 18:-?-:-?-.-?-- | 280.000 *
8 | Feb 23 2000 21:10:-?-.-?-- | Feb 24 2000 21:-?-:-?-.-?-- | 110.000 *

* The EXACT time of these burns has yet to be made available to the IOTs.

The last burn is intended as a “Trimming Burn” to be used to polish off the final orbital path.

During the burns, a short period of MODIS data loss occurred following the burn due to priority of Spacecraft vs. Instrument activities. This is not an anomaly, this is an operational agreement between the IOTs and FOT. The MODIS data losses occurred at:
043-20:43:27 – 043/22:15:29
049-22:45:29 – 049-23:26:01

Limited Life Item Status:

All limited life items are well within lifetime ranges. The precise statistics for each item have been received from LMMS/Valley Forge and will be incorporated next week.